# Isbell’s subfactor projections in a noetherian form

### v1

@@ 1 Introduction
  • The categories of distributive and modular lattices with modular connections as morphisms (a modular connection is a Galois connection satisfying the formulas given in Axiom 4 in this paper), because these categories are Grandis exact.
@@
In this paper we rely on the groundwork material pertaining to noetherian forms as laid down in [6]. For reader’s convenience, however, we nevertheless include below, in Section 2, a brief refresher on basic terminology, notation, axioms and results in the context of a noetherian form. We include in the same section some new observations that can be easily deduced from the results contained in [6] and which help with the rest of the paper. In Section 7, we extend Isbell’s work from [7] to the context of a noetherian form. In the last Section 8, we provide a counterexample to one of the claims in the refinement theorem formulated in [7].
2 Preliminaries
@@
A noetherian form is a form satisfying the axioms stated below. With the terminology recalled in the Introduction, in the context of a noetherian form $F$, the subgroups of a group $G$ form a poset — the fibre $F^{-1}(G)$, which we write as $\mathsf{Sub}G$. For any homomorphism $f\colon G\to H$, we write $X\subseteq_{f}Y$ where $X\in\mathsf{Sub}G$ and $Y\in\mathsf{Sub}H$, when there exists a morphism $f^{\prime}\colon X\to Y$ such that $F(f^{\prime})=f$. In the case when $f=\mathsf{id}_{G}$, this is the partial order relation of $\mathsf{Sub}G$, which we denote simply by $\subseteq$’ (drop the subscript). Note that the transitivity law of this partial order is a special case of a more general property, which arises from the fact that $F$ preserves composition:
@@
${\mathsf{Sub}G\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$${\mathsf{Sub}H\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f^{-1}}$
@@
hold for all $X\in\mathsf{Sub}G$ and $Y\in\mathsf{Sub}H$. Furthermore, for two composable homomorphisms $f\colon G\to H$ and $g\colon E\to G$, and subgroups $W$ of $E$ and $Y$ of $H$, we have $(fg)(W)=f(g(W))$ (and $(fg)^{-1}(Y)=g^{-1}(f^{-1}(Y))$, which can be derived). Lastly, for any group $G$ and subgroup $X$ of $G$, we have $\mathsf{id}_{G}(X)=X$ (and $\mathsf{id}_{G}^{-1}(X)=X$, which can be derived).
The axiom above implies that the two maps $f(-),f^{-1}(-)$ form a monotone Galois connection (i.e., an adjunction) between $\mathsf{Sub}X$ and $\mathsf{Sub}Y$, where the direct image map $f(-)$ is the left adjoint and the inverse image map $f^{-1}(-)$ is a right adjoint. Furthermore, it is easy to check that this Galois connection is in fact uniquely determined by $f$, thanks to the equivalences required in the axiom above. The usual and well-known properties of Galois connections (or more generally, of adjunctions) are sufficient to derive the conclusions claimed in brackets in the axiom above. The same usual properties yield that the direct image maps preserve joins (when they exist) and the inverse image maps preserve meets (when they exist).
For a group $G$, we denote the bottom and top elements of $\mathsf{Sub}G$, if they exist, by $0_{G}$ and $1_{G}$ respectively (the empty join and the empty meet), where the subscripts are dropped when it is clear which group we are referring to. Consider a homomorphism $f\colon G\to H$. The kernel of $f$, denoted $\mathsf{Ker}f$, is defined as $f^{-1}(0_{H})$ (the inverse image of the smallest subgroup of $H$). Similarly, the image of $f$, denoted $\mathsf{Im}f$, is defined as $f(1_{G})$ (the direct image of the largest subgroup). These terms obtain their usual meaning in usual group theory. The axiom above guarantees that similarly as in usual group theory, the kernel of an isomorphism is the least subgroup and the image is the largest subgroup.
@@
Axiom 4
For any group $G$, the subgroup poset $\mathsf{Sub}G$ is a bounded lattice. Moreover, for any homomorphism $f\colon G\rightarrow H$ and subgroups $X$ of $G$ and $Y$ of $H$, we have
@@
Theorem 4.1 (Lattice Isomorphism Theorem)
For any homomorphism $f\colon G\to H$, the direct and inverse image maps along $f$ result in an isomorphism of the intervals $[\mathsf{Ker}f,1_{G}]$ and $[0_{H},\mathsf{Im}f]$, which are sublattices of $\mathsf{Sub}G$ and $\mathsf{Sub}H$ (i.e., they are closed under binary meet and join), respectively.
See [9] for various other equivalent reformulations of the axiom above and links with closely related axioms considered in universal and categorical algebra. The simplest reformulation, which only makes use of the poset structure of lattices, is given by requiring that the set $\{f(X)\mid X\in\mathsf{Sub}G\}$ is down-closed and the set $\{f^{-1}(Y)\mid Y\in\mathsf{Sub}H\}$ is up-closed. In [9], following the terminology of Grandis [4, 5], we may call monotone maps $f$ between posets that admit a right adjoint $f^{-1}$ such that the two maps have these properties modular maps. A useful fact about modular maps is that their composite is also modular. In [4], a Galois connection between modular lattices is called a modular connection when its left adjoint is modular in our sense.
The second of the last two equalities in Axiom 4.1 is a weaker form of what, in the context of ‘hyperdoctrines’, Lawvere coined Frobenius reciprocity in [12]:
$f(f^{-1}(Y)\wedge X)=Y\wedge f(X).$ (4.2)
@@
$X\subseteq Z\quad\Rightarrow\quad(X\vee Y)\wedge Z=X\vee(Y\wedge Z).$ (4.3)
@@
Theorem 4.4 (Restricted Modular Law, [6])
For any three subgroups, $X$, $Y$, and $Z$ of a group $G$, if $Y$ is normal and $Z$ is conormal, then the implication (4.3) holds. The same implication holds also when $Y$ is conormal and $X$ is normal.
@@
$X\subseteq Z\quad\Rightarrow\quad X(Y\cap Z)=XY\cap Z.$ (4.5)
This implication holds even when $X$ and $Y$ are merely subsets of $G$. So, in particular, it holds when $X$ and $Y$ are subgroups of $G$ with $X$ not necessarily being a normal subgroup of $G$. In our abstract context, we cannot speak about subsets of groups unless they are subgroups, and neither can we consider the product of two subsets; the second half of the restricted modular law is then the closest we can get to (4.5).
Consider a subgroup $X$ of a group $G$. An embedding associated to $X$ is a homomorphism $\iota_{X}\colon X\to G$, which may or may not exist, defined by a universal property. Note that here the group $X$, i.e., the domain of $\iota_{X}$, is not the same thing as the subgroup $X$ (they are in fact objects in different categories), even though we write them the same way. In usual group theory, both $X$’s will will be the same group. The universal property states: $\mathsf{Im}\iota_{X}\subseteq X$ and for any homomorphism $f\colon H\to G$ with $\mathsf{Im}f\subseteq X$, there exists a unique homomorphism $f^{\prime}\colon H\to X$ which satisfies $f=\iota_{X}f^{\prime}$. It is easy to see that the identity homomorphism $\mathsf{id}_{G}\colon G\to G$ is an embedding associated to the largest subgroup $1_{G}$ of a group $G$. It is then in line with the convention above to write $G$ for the largest subgroup $1_{G}$ of $G$.
@@
Another useful property of embeddings and projection homomorphisms that was noted in [6] is that any embedding is an an embedding associated to its image, and moreover, if an embedding is associated with a conormal subgroup than its image equals that conormal subgroup. Dually, any projection homomorphism is a projection homomorphism associated with its kernel and if it is associated with a normal subgroup, than the kernel equals that normal subgroup.
@@
Theorem 6.1 (Homomorphism Induction Theorem, [6])
For a zigzag to induce a homomorphism it is necessary and sufficient that chasing the least subgroup of the initial node forward along the zigzag results in the least subgroup of the final node, and chasing the largest subogroup of the final node backward along the zigzag results in the largest subgroup of the initial node. Moreover, when a zigzag induces a homomorphism, the induced homomorphism is unique and the direct and inverse image maps of the induced homomorphism are given by chasing a subgroup forward and backward, respectively, along the zigzag.
@@
For two subgroups $X\lhd Y$ of a group $G$, chasing the interval $[0,1]=[0_{Y/X},Y/X]$ in $\mathsf{Sub}(Y/X)$ backward along the zigzag (6.2) results in the interval $[X,Y]$. Furthermore, chasing subgroups backward along the zigzag (6.2) gives an isomorphism between the lattice $\mathsf{Sub}(Y/X)$ and the sublattice $[X,Y]$ of $G$, whose inverse is given by chasing a subgroup $Z\in[X,Y]$ forward along the same zigzag. Finally, for any subgroup $Z$ of $G$, chasing $Z$ forward and then backward along the zigzag (6.2) results in the subgroup $(Z\wedge Y)\vee X$.
Proof 6.3.
The effect of chasing $Z$ forward and then backward along the zigzag (6.2) is:
@@
Since $\iota_{Y}$ is an embedding and $\pi_{\iota^{-1}_{Y}(X)}$ is a projection, chasing a subgroup backward along the zigzag (6.2) and then forward gives back the same subgroup. So we get an isomorphism between $\mathsf{Sub}(Y/X)$ and the subposet of $\mathsf{Sub}(G)$ given by those $Z\in\mathsf{Sub}(G)$ such that $Z=(Z\wedge Y)\vee X$. It is easy to see that the sublattice $[X,Y]$ of $\mathsf{Sub}(G)$ is in fact such subposet of $\mathsf{Sub}(G)$.
@@
For any three subgroups, $X$, $Y$, and $Z$ of a group $G$, if for some subgroup $S$ of $G$ we have $Y\lhd S$, $Z\subseteq S$ and $Z$ is conormal, then the implication (4.3) holds. The same implication holds when $Y$ is conormal, $Y\subseteq S$, $Z\subseteq S$ and $X\lhd S$ for some subgroup $S$ of $G$.
Proof 6.7.
@@
Expression Dual Expression
$G\in\mathbb{C}$ $G\in\mathbb{C}$
$X\in F^{-1}(G)$ $X\in F^{-1}(G)$
$X\subseteq_{f}Y$ $Y\subseteq_{f}X$
$gf$ $fg$
Table 1: Scheme for functorial duality of the internal language of a form $F\colon\mathbb{X}\to\mathbb{C}$
7 Isbell’s results on subfactors
In what follows we work in $\mathsf{Sub}G$, for a fixed $G$. An interval $X=[X^{-},X^{+}]$ is a subfactor (of $G$) when $X^{-}\lhd X^{+}$. The projection of a subgroup $Z$ in an interval $X=[X^{-},X^{+}]$ is defined as
@@ 7 Isbell’s results on subfactors
Proof 7.3.
By Lemma 6.2, $XY=Y$ is equivalent to the interval $[0,1]$ at $X^{+}/X^{-}$ chasing to the interval $[0,1]$ at $Y^{+}/Y^{-}$. Symmetrically, $YX=X$ is equivalent the interval $[0,1]$ at $Y^{+}/Y^{-}$ chasing to the interval $[0,1]$ at $X^{+}/X^{-}$. So the result follows from the Universal Isomorphism Theorem.
@@
$(YX)^{-}\wedge(XY)^{+}\subseteq(YX)^{-}\wedge Y^{+}\subseteq((YX)^{-}\wedge Y^{+})\vee Y^{-}=(YX)^{-}Y=((YX)Y)^{-}=(XY)^{-}.$
@@
$(YX)^{+}\wedge(XY)^{+}\subseteq(YX)^{+}\wedge Y^{+}\subseteq((YX)^{+}\wedge Y^{+})\vee Y^{-}=(YX)^{+}Y=((YX)Y)^{+}=(XY)^{+}.$
So $(YX)(XY)=XY$. Symmetrically, we get $(XY)(YX)=YX$. By Lemma 7.1, when $Y^{+}X$ and $X^{+}Y$ are conormal, $XY$ and $YX$ are subfactors. Since they project onto each other by the first part of the lemma, the corresponding quotients $Y^{+}X/Y^{-}X=(YX)^{+}/(YX)^{-}$ and $X^{+}Y/X^{-}Y=(XY)^{+}/(XY)^{-}$ must then be isomorphic by Theorem 7.1.
@@
$G=X_{0}\supset X_{1}\supset\dots\supset X_{n}=0_{G}$ (7.2)
such that each interval $[X_{i+1},X_{i}]$ is a subfactor. Here $A\supset B$ is the conjunction of $B\subseteq A$ and $A\neq B$. A refinement of a subnormal series is a subnormal series that contains all the subgroups of the original series. Following Isbell, we further define a subnormal series (7.2) of a group $G$ to be projectively isomorphic to a subnormal series
$G=Y_{0}\supset Y_{1}\supset\dots\supset Y_{m}=0_{G}$ (7.3)
@@
Given a subnormal series (7.3), if $[Y_{j+1},Y_{j}]$ projects onto a subfactor $X=[X^{-},X^{+}]$, then $X^{-}$ and $X^{+}$ have the same projection in any other subfactor $[Y_{k+1},Y_{k}]$ where $k\neq j$. In particular, if a subnormal series (7.2) is projectively isomorphic to a subnormal series (7.3), then the two subnormal series have the same number of terms and each subfactor $[X_{i+1},X_{i}]$ in the first series projects onto a unique subfactor $[Y_{j+1},Y_{j}]$ in the other series (and $[Y_{j+1},Y_{j}]$ projects onto $[X_{i+1},X_{i}]$).
Proof 7.7.
@@
For the second part, suppose a subnormal series (7.2) is projectively isomorphic to a subnormal series (7.3). Then each subfactor $[X_{i+1},X_{i}]$ in the first series projects onto a subfactor $[Y_{j+1},Y_{j}]$ of the second series such that $[Y_{j+1},Y_{j}]$ projects onto $[X_{i+1},X_{i}]$. Now, by the first part of the lemma, $X_{i+1},X_{i}$ will project to the same subgroup in any other subfactor $[Y_{k+1},Y_{k}]$, $k\neq j$ in (7.3). Since all subgroups in the series (7.3) are distinct, this implies that $[Y_{j+1},Y_{j}]$ is the unique subfactor in (7.3) onto which $[X_{i+1},X_{i}]$ projects. Next, we will show that for every subfactor $[Y_{j+1},Y_{j}]$ in (7.3) there is a unique subfactor $[X_{i+1},X_{i}]$ in (7.3) which projects onto $[Y_{j+1},Y_{j}]$. Project all subgroups from the series (7.2) in the subfactor $[Y_{j+1},Y_{j}]$. By monotonicity of projection, we will get a decreasing sequence of subgroups in the interval $[Y_{j+1},Y_{j}]$. This sequence will start with $Y_{j}$ (projection of $1$) and end with $Y_{j+1}$ (projection of 00). Recall that each $[X_{i+1},X_{i}]$ either projects onto $[Y_{j+1},Y_{j}]$ or else, $X_{i+1},X_{i}$ project to the same subgroup in $[Y_{j+1},Y_{j}]$. This implies that there must exist unique $[X_{i+1},X_{i}]$ which projects onto $[Y_{j+1},Y_{j}]$, which is what we wanted to show. Thus, the mapping a subfactor from (7.2) to a subfactor in (7.3) onto which it projects is a bijection from the set of subfactors $[X_{i+1},X_{i}]$ in (7.2) to the set of subfactors $[Y_{j+1},Y_{j}]$ in (7.3). This implies that the series (7.2) and (7.3) must have the same number of terms.
@@
Given two subnormal series (7.2) and (7.3) in a group $G$, the projections of the $Y_{j}$ in all $[X_{i+1},X_{i}]$ and the projections of the $X_{i}$ in all $[Y_{j+1},Y_{j}]$ form projectively isomorphic subnormal series, provided all of these projections except possibly $0_{G}$ are conormal.
Proof 7.8.
@@
$X_{0}=Y_{0}^{0}\supseteq\dots\supseteq Y_{m}^{0}=X_{1}=Y_{0}^{1}\supseteq\dots\supseteq Y_{m}^{1}=X_{2}=Y_{0}^{2}\supseteq\dots\supseteq Y_{m}^{n-1}=X_{n}$ (7.5)
@@
$Y_{0}=X_{0}^{0}\supseteq\dots\supseteq X_{n}^{0}=Y_{1}=X_{0}^{1}\supseteq\dots\supseteq X_{n}^{1}=Y_{2}=X_{0}^{2}\supseteq\dots\supseteq X_{n}^{m-1}=Y_{m}$ (7.6)
in the second case. Note that these two series have the same number of terms: there are $mn+1$ many terms in each case. If all terms in these series are conormal (except possibly 00), then each $[Y_{j+1}^{i},Y_{j}^{i}]$ and each $[X_{i+1}^{j},X_{i}^{j}]$ is a subfactor, by Lemma 7.4. Thus, eliminating duplicate terms in (7.5) and in (7.6) results in subnormal series. Consider a subfactor $[Y_{j+1}^{i},Y_{j}^{i}]$ from the first subnormal series. Then $Y_{j+1}^{i}\neq Y_{j}^{i}$. Since

### v2

@@ 1. Introduction
  • The categories of distributive and modular lattices with modular connections as morphisms (a modular connection is a Galois connection satisfying the formulas given in Axiom 2.2 in this paper), because these categories are Grandis exact.
@@
In this paper we rely on the groundwork material pertaining to noetherian forms as laid down in [8]. For reader’s convenience, however, we nevertheless include below, in Section 2, a brief refresher on basic terminology, notation, axioms and results in the context of a noetherian form. We include in the same section some new observations that can be easily deduced from the results contained in [8] and which help with the rest of the paper. In Section 3, we extend Isbell’s work from [12] to the context of a noetherian form. In Section 4, we provide a counterexample to one of the claims in the refinement theorem formulated in [12]. In Section 5, we briefly discuss some similarities and differences with other abstract approaches to the butterfly lemma and the refinement theorem in the literature.
2. Preliminaries
@@
A noetherian form is a form satisfying the axioms stated below. With the terminology recalled in the Introduction, in the context of a noetherian form $F$, the subgroups of a group $G$ form a poset — the fibre $F^{-1}(G)$, which we write as $\mathsf{Sub}(G)$. For any homomorphism $f\colon G\to H$, we write $X\subseteq_{f}Y$ where $X\in\mathsf{Sub}(G)$ and $Y\in\mathsf{Sub}(H)$, when there exists a morphism $f^{\prime}\colon X\to Y$ such that $F(f^{\prime})=f$. In the case when $f=\mathsf{id}_{G}$, this is the partial order relation of $\mathsf{Sub}(G)$, which we denote simply by $\subseteq$’ (drop the subscript). Note that the transitivity law of this partial order is a special case of a more general property, which arises from the fact that $F$ preserves composition:
@@
${\mathsf{Sub}(G)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$${\mathsf{Sub}(H)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f^{-1}}$
@@
hold for all $X\in\mathsf{Sub}(G)$ and $Y\in\mathsf{Sub}(H)$. Furthermore, for two composable homomorphisms $f\colon G\to H$ and $g\colon E\to G$, and subgroups $W$ of $E$ and $Y$ of $H$, we have $(fg)(W)=f(g(W))$ (and $(fg)^{-1}(Y)=g^{-1}(f^{-1}(Y))$, which can be derived). Lastly, for any group $G$ and subgroup $X$ of $G$, we have $\mathsf{id}_{G}(X)=X$ (and $\mathsf{id}_{G}^{-1}(X)=X$, which can be derived).
The axiom above implies that the two maps $f(-),f^{-1}(-)$ form a monotone Galois connection (i.e., an adjunction) between $\mathsf{Sub}(G)$ and $\mathsf{Sub}(H)$, where the direct image map $f(-)$ is the left adjoint and the inverse image map $f^{-1}(-)$ is a right adjoint. Furthermore, it is easy to check that this Galois connection is in fact uniquely determined by $f$, thanks to the equivalences required in the axiom above. The usual and well-known properties of Galois connections (or more generally, of adjunctions) are sufficient to derive the conclusions claimed in brackets in the axiom above. The same usual properties yield that the direct image maps preserve joins (when they exist) and the inverse image maps preserve meets (when they exist).
For a group $G$, we denote the bottom and top elements of $\mathsf{Sub}(G)$, if they exist, by $0_{G}$ and $1_{G}$ respectively (the empty join and the empty meet), where the subscripts are dropped when it is clear which group we are referring to. Consider a homomorphism $f\colon G\to H$. The kernel of $f$, denoted $\mathsf{Ker}f$, is defined as $f^{-1}(0_{H})$ (the inverse image of the smallest subgroup of $H$). Similarly, the image of $f$, denoted $\mathsf{Im}f$, is defined as $f(1_{G})$ (the direct image of the largest subgroup). These terms obtain their usual meaning in usual group theory. The axiom above guarantees that similarly as in usual group theory, the kernel of an isomorphism is the least subgroup and the image is the largest subgroup.
@@
Expression Dual Expression
$G\in\mathbb{C}$ $G\in\mathbb{C}$
$X\in F^{-1}(G)$ $X\in F^{-1}(G)$
$X\subseteq_{f}Y$ $Y\subseteq_{f}X$
$gf$ $fg$
Table 1. Scheme for functorial duality of the internal language of a form $F\colon\mathbb{X}\to\mathbb{C}$
Axiom 2.2.
For any group $G$, the subgroup poset $\mathsf{Sub}(G)$ is a bounded lattice. Moreover, for any homomorphism $f\colon G\rightarrow H$ and subgroups $X$ of $G$ and $Y$ of $H$, we have
@@
Theorem 2.3 (Lattice Isomorphism Theorem).
For any homomorphism $f\colon G\to H$, the direct and inverse image maps along $f$ result in an isomorphism of the intervals $[\mathsf{Ker}f,1_{G}]$ and $[0_{H},\mathsf{Im}f]$, which are sublattices of $\mathsf{Sub}(G)$ and $\mathsf{Sub}(H)$ (i.e., they are closed under binary meet and join), respectively.
See [15] for various other equivalent reformulations of the axiom above and links with closely related axioms considered in universal and categorical algebra. The simplest reformulation, which only makes use of the poset structure of lattices, is given by requiring that the set $\{f(X)\mid X\in\mathsf{Sub}(G)\}$ is down-closed and the set $\{f^{-1}(Y)\mid Y\in\mathsf{Sub}(H)\}$ is up-closed. In [15], following the terminology of Grandis [10, 11], we may call monotone maps $f$ between posets that admit a right adjoint $f^{-1}$ such that the two maps have these properties modular maps. A useful fact about modular maps is that their composite is also modular. In [10], a Galois connection between modular lattices is called a modular connection when its left adjoint is modular in our sense.
The second of the last two equalities in Axiom 2.2 is a weaker form of what, in the context of ‘hyperdoctrines’, Lawvere coined Frobenius reciprocity in [18]:
(1) $f(f^{-1}(Y)\wedge X)=Y\wedge f(X).$
@@
(2) $X\subseteq Z\quad\Rightarrow\quad(X\vee Y)\wedge Z=X\vee(Y\wedge Z).$
@@
Theorem 2.4 (Restricted Modular Law, [8]).
For any three subgroups, $X$, $Y$, and $Z$ of a group $G$, if $Y$ is normal and $Z$ is conormal, then the implication (2) holds. The same implication holds also when $Y$ is conormal and $X$ is normal.
@@
(3) $X\subseteq Z\quad\Rightarrow\quad X(Y\cap Z)=XY\cap Z.$
This implication holds even when $X$ and $Y$ are merely subsets of $G$. So, in particular, it holds when $X$ and $Y$ are subgroups of $G$ with $X$ not necessarily being a normal subgroup of $G$. In our abstract context, we cannot speak about subsets of groups unless they are subgroups, and neither can we consider the product of two subsets; the second half of the restricted modular law is then the closest we can get to (3).
Consider a subgroup $X$ of a group $G$. An embedding associated to $X$ is a homomorphism $\iota_{X}\colon X\to G$, which may or may not exist, defined by a universal property. Note that here the group $X$, i.e., the domain of $\iota_{X}$, is not the same thing as the subgroup $X$ (they are in fact objects in different categories), even though we write them the same way. In usual group theory, both $X$’s will be the same group. The universal property states: $\mathsf{Im}\iota_{X}\subseteq X$ and for any homomorphism $f\colon H\to G$ with $\mathsf{Im}f\subseteq X$, there exists a unique homomorphism $f^{\prime}\colon H\to X$ which satisfies $f=\iota_{X}f^{\prime}$. It is easy to see that the identity homomorphism $\mathsf{id}_{G}\colon G\to G$ is an embedding associated to the largest subgroup $1_{G}$ of a group $G$. It is then in line with the convention above to write $G$ for the largest subgroup $1_{G}$ of $G$.
@@
Another useful property of embeddings and projection homomorphisms that was noted in [8] is that any embedding is an embedding associated to its image, and moreover, if an embedding is associated with a conormal subgroup, then its image equals that conormal subgroup. Dually, any projection homomorphism is a projection homomorphism associated with its kernel and if it is associated with a normal subgroup, then the kernel equals that normal subgroup.
@@
Theorem 2.7 (Homomorphism Induction Theorem, [8]).
For a zigzag to induce a homomorphism it is necessary and sufficient that chasing the least subgroup of the initial node forward along the zigzag results in the least subgroup of the final node, and chasing the largest subgroup of the final node backward along the zigzag results in the largest subgroup of the initial node. Moreover, when a zigzag induces a homomorphism, the induced homomorphism is unique and the direct and inverse image maps of the induced homomorphism are given by chasing a subgroup forward and backward, respectively, along the zigzag.
@@
For two subgroups $X\lhd Y$ of a group $G$, chasing the interval $[0,1]=[0_{Y/X},Y/X]$ in $\mathsf{Sub}(Y/X)$ backward along the zigzag (4) results in the interval $[X,Y]$. Furthermore, chasing subgroups backward along the zigzag (4) gives an isomorphism between the lattice $\mathsf{Sub}(Y/X)$ and the sublattice $[X,Y]$ of $G$, whose inverse is given by chasing a subgroup $Z\in[X,Y]$ forward along the same zigzag. Finally, for any subgroup $Z$ of $G$, chasing $Z$ forward and then backward along the zigzag (4) results in the subgroup $(Z\wedge Y)\vee X$.
Proof.
The effect of chasing $Z$ forward and then backward along the zigzag (4) is:
@@
Since $\iota_{Y}$ is an embedding and $\pi_{\iota^{-1}_{Y}(X)}$ is a projection, chasing a subgroup backward along the zigzag (4) and then forward gives back the same subgroup. So we get an isomorphism between $\mathsf{Sub}(Y/X)$ and the subposet of $\mathsf{Sub}(G)$ given by those $Z\in\mathsf{Sub}(G)$ such that $Z=(Z\wedge Y)\vee X$. It is easy to see that the sublattice $[X,Y]$ of $\mathsf{Sub}(G)$ is in fact such subposet of $\mathsf{Sub}(G)$. ∎
@@
For any three subgroups, $X$, $Y$, and $Z$ of a group $G$, if for some subgroup $S$ of $G$ we have $Y\lhd S$, $Z\subseteq S$ and $Z$ is conormal, then the implication (2) holds. The same implication holds when $Y$ is conormal, $Y\subseteq S$, $Z\subseteq S$ and $X\lhd S$ for some subgroup $S$ of $G$.
Proof.
@@
3. Isbell’s results on subfactors
In what follows we work in $\mathsf{Sub}(G)$, for a fixed $G$. An interval $X=[X^{-},X^{+}]$ is a subfactor (of $G$) when $X^{-}\lhd X^{+}$. The projection of a subgroup $Z$ in an interval $X=[X^{-},X^{+}]$ is defined as
@@ 3. Isbell’s results on subfactors
Proof.
By Lemma 2.9, $XY=Y$ is equivalent to the interval $[0,1]$ at $X^{+}/X^{-}$ chasing to the interval $[0,1]$ at $Y^{+}/Y^{-}$. Symmetrically, $YX=X$ is equivalent the interval $[0,1]$ at $Y^{+}/Y^{-}$ chasing to the interval $[0,1]$ at $X^{+}/X^{-}$. So the result follows from the Universal Isomorphism Theorem. ∎
@@
$(YX)^{-}\wedge(XY)^{+}$ $\subseteq(YX)^{-}\wedge Y^{+}$
$\subseteq((YX)^{-}\wedge Y^{+})\vee Y^{-}$
$=(YX)^{-}Y$
$=((YX)Y)^{-}$
$=(XY)^{-}.$
@@
$(YX)^{+}\wedge(XY)^{+}$ $\subseteq(YX)^{+}\wedge Y^{+}$
$\subseteq((YX)^{+}\wedge Y^{+})\vee Y^{-}=(YX)^{+}Y$
$=((YX)Y)^{+}$
$=(XY)^{+}.$
So $(YX)(XY)=XY$. Symmetrically, we get $(XY)(YX)=YX$. By Lemma 3.1, when $Y^{+}X$ and $X^{+}Y$ are conormal, $XY$ and $YX$ are subfactors. Since they project onto each other by the first part of the lemma, the corresponding quotients
$Y^{+}X/Y^{-}X=(YX)^{+}/(YX)^{-}\textrm{ and }X^{+}Y/X^{-}Y=(XY)^{+}/(XY)^{-}$
must then be isomorphic by Theorem 3.2. ∎
@@
(5) $G=X_{0}\supset X_{1}\supset\dots\supset X_{n}=0_{G}$
such that each interval $[X_{i+1},X_{i}]$ is a subfactor. Here $A\supset B$ is the conjunction of $B\subseteq A$ and $A\neq B$. A refinement of a subnormal series is a subnormal series that contains all the subgroups of the original series. Following Isbell, we further define a subnormal series (5) of a group $G$ to be projectively isomorphic to a subnormal series
(6) $G=Y_{0}\supset Y_{1}\supset\dots\supset Y_{m}=0_{G}$
@@
Given a subnormal series (6), if $[Y_{j+1},Y_{j}]$ projects onto a subfactor $X=[X^{-},X^{+}]$, then $X^{-}$ and $X^{+}$ have the same projection in any other subfactor $[Y_{k+1},Y_{k}]$ where $k\neq j$. In particular, if a subnormal series (5) is projectively isomorphic to a subnormal series (6), then the two subnormal series have the same number of terms and each subfactor $[X_{i+1},X_{i}]$ in the first series projects onto a unique subfactor $[Y_{j+1},Y_{j}]$ in the other series (and $[Y_{j+1},Y_{j}]$ projects onto $[X_{i+1},X_{i}]$).
Proof.
@@
For the second part, suppose a subnormal series (5) is projectively isomorphic to a subnormal series (6). Then each subfactor $[X_{i+1},X_{i}]$ in the first series projects onto a subfactor $[Y_{j+1},Y_{j}]$ of the second series such that $[Y_{j+1},Y_{j}]$ projects onto $[X_{i+1},X_{i}]$. Now, by the first part of the lemma, $X_{i+1}$ and $X_{i}$ will project to the same subgroup in any other subfactor $[Y_{k+1},Y_{k}]$, $k\neq j$ in (6). Since all subgroups in the series (6) are distinct, this implies that $[Y_{j+1},Y_{j}]$ is the unique subfactor in (6) onto which $[X_{i+1},X_{i}]$ projects. Next, we will show that for every subfactor $[Y_{j+1},Y_{j}]$ in (6) there is a unique subfactor $[X_{i+1},X_{i}]$ in (6) which projects onto $[Y_{j+1},Y_{j}]$. Project all subgroups from the series (5) in the subfactor $[Y_{j+1},Y_{j}]$. By monotonicity of projection, we will get a decreasing sequence of subgroups in the interval $[Y_{j+1},Y_{j}]$. This sequence will start with $Y_{j}$ (projection of $1$) and end with $Y_{j+1}$ (projection of 00). Recall that each $[X_{i+1},X_{i}]$ either projects onto $[Y_{j+1},Y_{j}]$ or else, $X_{i+1}$ and $X_{i}$ project to the same subgroup in $[Y_{j+1},Y_{j}]$. This implies that there must exist unique $[X_{i+1},X_{i}]$ which projects onto $[Y_{j+1},Y_{j}]$, which is what we wanted to show. Thus, the mapping a subfactor from (5) to a subfactor in (6) onto which it projects is a bijection from the set of subfactors $[X_{i+1},X_{i}]$ in (5) to the set of subfactors $[Y_{j+1},Y_{j}]$ in (6). This implies that the series (5) and (6) must have the same number of terms. ∎
@@
Given two subnormal series (5) and (6) in a group $G$, the projections of the $Y_{j}$ in all $[X_{i+1},X_{i}]$ and the projections of the $X_{i}$ in all $[Y_{j+1},Y_{j}]$ form projectively isomorphic subnormal series, provided all of these projections except possibly $0_{G}$ are conormal.
Proof.
@@
(7) $X_{0}=Y_{0}^{0}\supseteq\dots\supseteq Y_{m}^{0}=X_{1}=Y_{0}^{1}\supseteq\dots\supseteq Y_{m}^{1}=X_{2}=Y_{0}^{2}\supseteq\dots\supseteq Y_{m}^{n-1}=X_{n}$
@@
(8) $Y_{0}=X_{0}^{0}\supseteq\dots\supseteq X_{n}^{0}=Y_{1}=X_{0}^{1}\supseteq\dots\supseteq X_{n}^{1}=Y_{2}=X_{0}^{2}\supseteq\dots\supseteq X_{n}^{m-1}=Y_{m}$
in the second case. Note that these two series have the same number of terms: there are $mn+1$ many terms in each case. If all terms in these series are conormal (except possibly 00), then each $[Y_{j+1}^{i},Y_{j}^{i}]$ and each $[X_{i+1}^{j},X_{i}^{j}]$ is a subfactor, by Lemma 3.3. Thus, eliminating duplicate terms in (7) and in (8) results in subnormal series. Consider a subfactor $[Y_{j+1}^{i},Y_{j}^{i}]$ from the first subnormal series. Then $Y_{j+1}^{i}\neq Y_{j}^{i}$. Since